why is 3 & 2/3 and 5 & 1/4 equal 77/4?
Please respond A.S.A.P
I NEED ANSWER BY TOMORROW MORNING AT 7:00
IF NOT I WILL BE DISAPPOINTED :)

Answer by oscargut(2103)

3 and 2/3 = 9/3+2/3 = 11/3
5 and 1/4 = 5+ 1/4 = 21/4
(11/3)(21/4) = (11*21)/(3*4) = (11*7)/4 = 77/4
If you need more help with this kind if questions, you can contact me now at:
mthman@gmail.com
Thanks